# Depth Restoration of Hand-Held Transparent Objects for Human-to-Robot Handover

### v1

@@ Abstract
Transparent objects are common in daily life, while their unique optical properties pose challenges for RGB-D cameras, which struggle to capture accurate depth information. For assistant robots, accurately perceiving transparent objects held by humans is essential for effective human-robot interaction. This paper presents a Hand-Aware Depth Restoration (HADR) method for hand-held transparent objects based on creating an implicit neural representation function from a single RGB-D image. The proposed method introduces the hand posture as an important guidance to leverage semantic and geometric information. To train and evaluate the proposed method, we create a high-fidelity synthetic dataset called TransHand-14K with a real-to-sim data generation scheme. Experiments show that our method has a better performance and generalization ability compared with existing methods. We further develop a real-world human-to-robot handover system based on the proposed depth restoration method, demonstrating its application value in human-robot interaction.
@@ I Introduction
RGB-D cameras have become popular for robotics perception due to their ability to capture 3D information, which is crucial for robot grasping [1, 2] and manipulation [3, 4]. However, transparent objects leave challenges in capturing correct depth data from RGB-D cameras due to refraction and reflection of light [5]. This issue is particularly critical for service robots [6], which are deployed in scenarios where transparent objects widely exist. Given the growing demand for human-robot interaction, it’s essential for robots to accurately perceive transparent objects held by humans and manipulate them safely. This requires enhanced perception accuracy and robustness, especially when dealing with hand occlusions.
Many attempts have been made to address the problem of transparent object depth estimation by introducing a depth restoration model [5, 7, 8]. The general idea behind those methods is extracting information from the RGB image and the corrupted depth image to predict the correct depth value at transparent objects’ regions. However, restoring the hand-held transparent objects presents many additional challenges and a research gap still exists. First, generating data for model training could be much harder in both real-world and simulation settings. Second, hand occlusions bring difficulties in depth restoration. Since objects’ appearances are highly related to the hand geometrically, the same object may look different because of hand occlusion, which leaves a higher demand for model generalization ability. Additionally, previous works on human-to-robot handover [9, 10] claim their methods are struggled with transparent objects. It remains uncertain whether introducing a depth restoration model can improve the overall performance of the transparent object handover.
@@
In this paper, we aim to solve the aforementioned problems from the perspectives of data, algorithm, and application. A synthetic dataset named TransHand-14K is first proposed, which is used for hand-held transparent object perception research. We introduce a real-to-sim data generation scheme to ensure the quality and fidelity of data. Based on the dataset TransHand-14K, we present a hand-aware depth restoration method for hand-held transparent objects. This method incorporates hand-holding postures to facilitate object reconstruction. Results demonstrate that our method surpasses existing methods, especially in terms of object-level generalization ability. Finally, we introduce a human-to-robot handover pipeline for transparent objects based on the proposed depth restoration method. This pipeline is adaptable to diverse commonly encountered transparent objects and can respond dynamically to human movements during the handover process.
In summary, the key contribution of our work includes the following three aspects: 1) a synthetic dataset TransHand-14K for transparent object perception research, 2) a novel hand-aware depth restoration method for hand-held transparent objects, 3) a human-to-robot transparent objects handover system deployed in the real world.
@@ II Related Work
II-A Depth Restoration for Transparent Objects
To minimize errors in the depth estimation of transparent objects using 3D sensors, various depth restoration methods are proposed, which utilize RGB and corrupted depth data to reconstruct the objects’ shape. Some early works restore the depth with multiple camera viewpoints based on visual hull [11], stereo-view matching [12], and NeRF [13, 14]. While multi-view methods suffer from relatively long inference time, other works reconstruct transparent objects only using single-view. [5] introduces a global optimization method that restores corrupted depth by predicting objects’ surface normals and edges. In [15], a local implicit neural representation method is proposed. [16] designs an end-to-end depth completion network with Unet architecture called DFNet. To better extract features from input data, transformer architecture [17] is applied to some recent works such as TODE-trans [7] and SwinDRNet [8]. However, all aforementioned methods are designed to reconstruct transparent objects on the tabletop. It is still an open problem to deal with hand-held transparent objects, which is a common scenario of robot and human interaction.
@@ II-B Datasets for Transparent Objects Perception
Training a depth restoration model for transparent objects requires both corrupted and perfect depth data from RGB-D cameras. These data can be generated from either the real world or the synthetic. [5] constructs a large-scale dataset with both synthetic data rendering with Blender and real-world data by spraying opaque material on transparent objects. In the work of [15], a synthetic dataset of transparent objects is generated by using the NVIDIA Omniverse platform [18]. Another large-scale real-world dataset called TransCG is introduced in [16], which generates data with an auto-collection pipeline. However, no dataset is currently available for grasped or hand-held transparent objects, as creating such a real-world dataset could require significant time and financial resources. Creating a synthetic dataset also presents challenges, particularly in generating realistic hand poses for objects of various shapes, which is crucial for narrowing the sim-to-real gap.
@@ III-A TransHand-14K Dataset Generation
Figure 2: (a) A real-to-sim data generation scheme. (b) We use MANUS VR glove to capture the pose data. (c) Eight transparent objects are included in the dataset.
@@
In this work, we propose a synthetic dataset specifically designed for hand-held transparent objects, named TransHand-14K. Generating such a dataset is more challenging than the conventional object-only datasets [5, 15]. Beyond considerations of background setting and lighting, it is also essential to account for the various grasping postures and their realism. Existing methods of hand pose generation method have some limitations [19]. For instance, the hand-held object dataset proposed by [20] only provides six types of grasping schemes for different objects. Additionally, the generated data suffers from problems of unrealistic background and hand poses.
@@
To solve those problems, we propose a real-to-sim data generation scheme. As shown in Fig. 2(a), we first utilize a MANUS VR glove [21] (Prime X Haptic) to capture the hand pose data of holding transparent objects, and then map the twin hand poses in Blender [22]. The synthetic hand in Blender is built on a skeleton model as shown in Fig. 2(b). To reduce the difficulty of modeling in Blender, we adopt a 22-point skeleton model by creating a virtual point (number 2), which is different from the commonly used 21-point skeleton model. The skeletal control can be achieved by feeding joint angle data captured by the MANUS VR glove. After generating the grasping poses, we manually place the corresponding transparent object into the proper position. 3D scanning is used to model eight common transparent objects with different shapes into Blender for rendering. In addition to ensuring the realism of generated hand posture, the proposed real-to-sim data generation method can easily obtain various grasping poses of the same object, significantly reducing the difficulty of dataset construction.
Figure 3: Visualization of proposed dataset TransHand-14K, which can be used for various hand-held transparent objects perception research.
@@
Figure 4: Overview of proposed depth restoration method. (a) Ray and voxel features are generated from the RGB and corrupted point cloud. (b) We introduce hand pose as an additional guidance for depth restoration. (c) The terminated probability and position of each ray-voxel pair are predicted. The final restored depth is obtained by ray-wise maxpooling.
@@
For the scene and background setting, we follow the work of Trans6D-32K [23]. The proposed dataset contains 14100 images with a resolution of 1280 $\times$ 720. As shown in Fig. 3, the proposed dataset contains labels of the hand 21-keypoint, objects mask, hand mask, and perfect depth image. In addition to depth completion, TransHand-14K can be used for other transparent object perception tasks such as detection, recognition, segmentation, and 6D pose estimation.
III-B Hand Aware Depth Restoration Method
@@
Ray-Voxel Implicit Neural Representation. We propose a depth restoration method for hand-held transparent objects. Inspired by the work of [15], the proposed method is based on generating a local implicit neural representation of ray-voxel pairs. The corrupted depth image is first transformed into the point cloud and placed into a voxel grid space in the camera coordinate. A pixel $I_{i}$ in the image can be viewed as a camera ray $r_{i}$ which passes through a set of occupied voxels $\mathbf{V}^{occ}$. The ray-voxel pair [15] is defined as $\phi_{ij}=(r_{i},v_{j})$, where $r_{i}$ is the camera ray direction for $i$-th pixel and $v_{j}$ is the intersected voxel with index $j$.
@@
$d_{i}=d_{ij^{\prime}}^{in}+\delta_{ij^{\prime}}r_{i},\>\>j^{\prime}=\mathop{\mathrm{argmax}}_{j:v_{j}^{\text{occ}}\in\mathbf{V}_{r_{i}}^{\text{occ}}}p_{ij}^{\text{end}},$ (1)
@@
where $d_{i}$ is the predicted depth of ray $r_{i}$; $d_{ij^{\prime}}^{in}$ is the entering position of ray-voxel intersected point; $\delta_{ij^{\prime}}$ is the offset distance; $p_{ij}^{\text{end}}$ is the probability the ray $r_{i}$ terminating in the voxel $v_{j}$. This expression shows that we can generate an implicit function based on neural networks to predict each ray-voxel pair’s terminated position and probability to realize the goal of restoring transparent objects’ depth.
@@
Ray and Voxel Feature Extraction. Given a ray-voxel pair $\phi_{ij}=(r_{i},v_{j})$, we have to generate its corresponding feature. Fig. 4(a) shows the model architecture of extracting ray-wise and voxel-wise features from the input RGB image and corrupted point cloud.
@@
For the input RGB image $I_{C}$, we use the swin transformer [24] as the backbone to extract multi-scale image features $\{F^{i}_{\text{rgb}}\}_{i=1,2,3,4}$, where $F^{i}_{\text{rgb}}\in\mathbb{R}^{\frac{H}{4i}\times\frac{W}{4i}\times iC}$ and $C=96$ is the embedding dimension of first layer output. We then introduce a features aggregation module to fuse multi-scale features into a single dense color feature map $F^{\text{dense}}_{\text{rgb}}\in\mathbb{R}^{H\times W\times C_{\text{rgb}}}$, where $C_{rgb}=32$. This is realized by first aligning different features into the same dimension of $\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times C}$, and concatenating them together. We then reduce the channels into $C_{\text{rgb}}$ using a $1\times 1$ convolution layer and conduct bilinear interpolation to resize the feature into $H\times W$. For the ray $r_{i}$, its corresponding ray feature can be extracted from $F^{\text{dense}}_{\text{rgb}}$ using ROI alignment [25], where the input size is $8\times 8$ and the output size is $2\times 2$. The final ray feature $F_{\text{ray}}$ is the flattened vector of the ROI alignment output.
@@
For the voxel feature, early work [15] generates it by sending the position and color of the valid point cloud into a voxel-based PointNet [26, 27]. However, this method could make it hard to fully leverage the semantic information of RGB input. To overcome this limitation, we introduce an early-stage point fusion method. We first encode the geometry embedding of point cloud $P_{\text{xyz}}\in\mathbb{R}^{N\times 16}$ using the point relative position to its voxel center. We then project the valid point cloud with camera intrinsic parameters onto the generated dense color feature map $F^{\text{dense}}_{\text{rgb}}$ and generate point-wise color embedding $P_{\text{rgb}}\in\mathbb{R}^{N\times 16}$ by the channel reduction. These two features are concatenated into the fused point feature $P_{\text{fuse}}$. The final voxel feature $F_{\text{voxel}}$ is generated from $P_{\text{fuse}}$ using a voxel feature encoder, which contains a set of voxel-wise maxpooling layers and fully connected layers [28].
@@
Hand Feature Extraction. In addition to the ray and voxel feature, we introduce hand pose as another important guidance for depth restoration. One of our insights is that the hand pose can provide both geometric and semantic context, facilitating the object reconstruction result. As shown in Fig. 4(b), we first extract the 21 hand keypoints by using the hand pose detector. In our experiment, a ResNet-18 [29] is adopted as the backbone to extract the hand feature map from the corrupted depth image, which is then transferred into 21 hand keypoints $X_{\text{hand}}\in\mathbb{R}^{21\times 3}$ through the weighted average regression [30].
@@
Two types of hand features are generated based on the detected hand keypoints. First, the keypoints are transferred into the hand-wrist coordinate and formed absolute feature $F_{\text{hand}}^{\text{abs}}$. It represents the absolute hand-holding pose, which is invariant to the change of camera position and direction. The second type of hand feature $F_{\text{hand}}^{\text{rel}}$ is the relative position between each hand keypoint and the center of intersected voxel $v_{j}$. This feature represents the spatial position relationship between occupied voxels and the hand. Absolute and relative hand features are flattened and contacted into the final hand feature $F_{\text{hand}}$.
@@
where $\oplus$ represents the feature-wise concatenation operator, $\gamma(r_{i})$ and $\gamma(d_{ij})$ are the ray direction and voxel position embedding, which has shown to be important for the final accuracy [15]. Two MLPs are used to decode the feature into the predicted terminated position and the existing probability of each ray-voxel pair. We obtain the final predicted depth value by using ray-wise maxing pooling [15], as expressed in Equation 1.
@@
III-C H2R Handover Strategy of Transparent Objects
Figure 5: Overview of the proposed handover strategy. The whole handover process is divided into three stages: Wait & Observe, Approach & React, Grasp & Retrieve.
In this section, we develop a handover strategy for transparent objects based on the proposed hand-aware depth restoration method. As shown in Fig. 5, we divide the whole human-to-robot handover process into three stages: 1) Wait & Observe, 2) Approach & React, 3) Grasp & Retrieve.

### v2

@@ Abstract
Transparent objects are common in daily life, while their optical properties pose challenges for RGB-D cameras to capture accurate depth information. This issue is further amplified when these objects are hand-held, as hand occlusions further complicate depth estimation. For assistant robots, however, accurately perceiving hand-held transparent objects is critical to effective human-robot interaction. This paper presents a Hand-Aware Depth Restoration (HADR) method based on creating an implicit neural representation function from a single RGB-D image. The proposed method utilizes hand posture as an important guidance to leverage semantic and geometric information of hand-object interaction. To train and evaluate the proposed method, we create a high-fidelity synthetic dataset named TransHand-14K with a real-to-sim data generation scheme. Experiments show that our method has better performance and generalization ability compared with existing methods. We further develop a real-world human-to-robot handover system based on HADR, demonstrating its potential in human-robot interaction applications.
@@ I Introduction
RGB-D cameras have become popular for robotic perception due to their ability to capture 3D information, which is crucial for robot grasping [1] and manipulation [2]. However, transparent objects pose challenges in capturing correct depth information from RGB-D cameras due to light refraction and reflection [3]. This issue is particularly critical for service robots [4], which are deployed in scenarios where transparent objects widely exist. Given the growing demand for human-robot interaction, it’s essential for robots to accurately perceive transparent objects held by humans and manipulate them safely. This further increases the need for precise and robust depth estimation of transparent objects.
Many attempts have been made to address the problem of transparent object depth estimation by introducing depth restoration models [3, 5, 6]. The general idea behind those methods is extracting information from the RGB and corrupted depth images to predict the correct depth value of transparent objects. However, restoring the hand-held transparent objects presents many new challenges and a research gap still exists. First, generating data for model training could be much harder in both real-world and simulation settings. At present, there exists no dataset supporting this area of research. Second, hand occlusions bring difficulties in depth restoration. Since objects’ appearances are highly related to the hand geometrically, the same object may look different because of hand occlusions, which leaves a higher demand for model generalization ability. Additionally, previous studies on human-to-robot (H2R) handover [7, 8] claim their methods are struggled with transparent objects. It remains uncertain whether incorporating depth restoration can improve the overall performance of the transparent object handover.
@@
In this paper, we aim to solve the aforementioned problems from the perspectives of data, algorithm, and application. A synthetic dataset named TransHand-14K is first proposed, which is used for hand-held transparent object perception research. We implement a real-to-sim data generation scheme to ensure the quality and fidelity of data. Based on the dataset TransHand-14K, we present a Hand-Aware Depth Restoration (HADR) method for hand-held transparent objects. This method incorporates hand-holding postures to facilitate object reconstruction. Results demonstrate that our method surpasses existing methods, especially in terms of object-level generalization ability. Furthermore, we introduce a human-to-robot handover workflow for transparent objects based on the proposed depth restoration method. This workflow is adaptable to a wide range of commonly encountered transparent objects and is capable of dynamically responding to human movements during the handover process.
In summary, the key contribution of our work includes the following three aspects: 1) a synthetic dataset TransHand-14K for transparent object perception research, 2) a novel depth restoration method HADR for hand-held transparent objects, 3) a human-to-robot transparent objects handover workflow deployed in a real-world system.
@@ II Related Work
II-A Depth Restoration of Transparent Objects
To minimize errors in the depth estimation of transparent objects using 3D sensors, various depth restoration methods are proposed, which utilize RGB and corrupted depth data to reconstruct the objects’ shape. Some early works restore the depth with multiple camera viewpoints based on visual hull [9], stereo-view matching [10], and NeRF [11, 12]. While multi-view methods suffer from relatively long inference time, other works reconstruct transparent objects only using single-view. A global optimization method is introduced in [3] that restores corrupted depth by predicting objects’ surface normals and edges. In [13], a local implicit neural representation method is proposed. Authors in [14] design an end-to-end depth completion network with Unet architecture called DFNet. To better extract features from input data, transformer architecture [15] is applied to some recent works such as TODE-trans [5] and SwinDRNet [6]. However, all aforementioned methods are designed to reconstruct transparent objects on the tabletop. Depth restoration of hand-held transparent objects remains an open problem, which is crucial for human-robot interaction.
@@ II-B Datasets for Transparent Objects Perception
Training a depth restoration model for transparent objects requires both corrupted and perfect depth data from RGB-D cameras. These data can be generated from either the real world or the simulation environments. A large-scale dataset is proposed in [3], which contains both synthetic data rendering with Blender and real-world data by spraying opaque material on transparent objects. In the work of [13], a synthetic dataset of transparent objects is generated by using the NVIDIA Omniverse platform [16]. Another large-scale real-world dataset called TransCG is introduced in [14], which generates data with an auto-collection pipeline. However, no dataset is currently available for grasped or hand-held transparent objects, as creating such a real-world dataset could require significant time and financial resources. Creating a synthetic dataset also presents challenges, particularly in generating realistic hand poses for objects with various shapes, which is crucial for narrowing the sim-to-real gap.
@@ III-A TransHand-14K Dataset Generation
Figure 2: TransHand-14K data generation method. (a) A real-to-sim hand pose generation scheme is introduced. (b) We use a MANUS VR glove to capture the pose data. (c) Eight transparent objects are included in TransHand-14K.
@@
In this work, we propose a synthetic dataset specifically designed for hand-held transparent objects, named TransHand-14K. Generating such a dataset is more challenging than the conventional object-only datasets [3, 13]. Beyond considerations of background setting and lighting, it is also essential to account for the various handhold postures and their realism. Existing works of hand pose generation have some limitations [17]. For instance, the hand-held object dataset proposed by [18] only provides six types of grasping schemes for different objects. Additionally, the generated data suffers from problems of unrealistic background and hand poses.
@@
To solve those problems, we propose a real-to-sim data generation scheme. As shown in Fig. 2(a), we first utilize a MANUS VR glove [19] (Prime X Haptic) to capture the hand pose data of holding transparent objects, and then map the twin hand poses in Blender [20]. The synthetic hand in Blender is built on a skeleton model as shown in Fig. 2(b). To reduce the difficulty of modeling in Blender, we adopt a 22-point skeleton model by creating a virtual point (number 2), which is different from the commonly used 21-point skeleton model. The skeletal control can be achieved by feeding joint angle data captured by the MANUS VR glove. After generating the handhold poses, we manually place the corresponding transparent object into the proper position. 3D scanning is used to model eight common transparent objects with different shapes into Blender for rendering. In addition to ensuring the realism of generated hand posture, the proposed real-to-sim data generation method can easily obtain various grasping poses of the same object, significantly reducing the difficulty of dataset construction.
Figure 3: Visualization of proposed dataset TransHand-14K.
@@
Figure 4: Overview of HADR. (a) Ray and voxel features are generated from the RGB and corrupted point cloud. (b) We introduce the handhold pose as an important guidance for depth restoration. This feature is related to geometric and semantic information of hand-object interaction. (c) The terminated probability and position of each ray-voxel pair are predicted. The final restored depth is obtained by ray-wise maxpooling.
@@
For the scene and background setting, we follow the work of Trans6D-32K [21]. Our dataset contains 14100 images with a resolution of 1280 $\times$ 720. As shown in Fig. 3, TransHand-14K contains RGB image, perfect depth image, objects mask, hand mask, and labels of the hand keypoints. In addition to depth completion, this dataset can be used for other transparent object perception tasks such as detection, recognition, segmentation, and 6D pose estimation.
III-B Hand-Aware Depth Restoration
@@
Ray-Voxel Implicit Neural Representation. We propose a depth restoration method named HADR for hand-held transparent objects. Inspired by the work of [13], our method is based on generating a local implicit neural representation of ray-voxel pairs. The corrupted depth image is first transformed into the point cloud and placed into a voxel grid space in the camera coordinate. A pixel $I_{i}$ in the image can be viewed as a camera ray $r_{i}$ which passes through a set of occupied voxels $\mathbf{V}^{occ}$. The ray-voxel pair [13] is defined as $\phi_{ij}=(r_{i},v_{j})$, where $r_{i}$ is the camera ray direction for $i$-th pixel and $v_{j}$ is the intersected voxel with index $j$.
@@
$d_{i}=d_{ij^{\prime}}^{\text{in}}+\delta_{ij^{\prime}}r_{i},\>\>j^{\prime}=\mathop{\mathrm{argmax}}_{j:v_{j}^{\text{occ}}\in\mathbf{V}_{r_{i}}^{\text{occ}}}p_{ij}^{\text{end}},$ (1)
@@
where $d_{i}$ is the predicted depth of ray $r_{i}$; $d_{ij^{\prime}}^{\text{in}}$ is the entering position of ray-voxel intersected point; $\delta_{ij^{\prime}}$ is the offset distance; $p_{ij}^{\text{end}}$ is the probability of ray $r_{i}$ terminated in the voxel $v_{j}$. This expression indicates that we can use neural networks to generate an implicit function for predicting the terminated position and probability of each ray-voxel pair, thereby restoring the accurate depth of transparent objects.
@@
Ray and Voxel Feature Extraction. To construct the implicit neural representation, we first generate ray-wise and voxel-wise features for each ray-voxel pair. Fig. 4(a) shows the model architecture of extracting these features from the input RGB image and corrupted point cloud.
@@
For the input RGB image $I_{C}$, we use the swin transformer [22] as the backbone to extract multi-scale image features $\{F^{i}_{\text{rgb}}\}_{i=1,2,3,4}$, where $F^{i}_{\text{rgb}}\in\mathbb{R}^{\frac{H}{4i}\times\frac{W}{4i}\times iC}$ and $C=96$ is the embedding dimension of first layer output. We then introduce a features aggregation module to fuse multi-scale features into a single dense color feature map $F^{\text{dense}}_{\text{rgb}}\in\mathbb{R}^{H\times W\times C_{\text{rgb}}}$, where $C_{\text{rgb}}=32$. This is realized by first aligning different features into the same dimension of $\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times C}$, and concatenating them together. We then reduce the channels into $C_{\text{rgb}}$ using a $1\times 1$ convolution layer and conduct bilinear interpolation to resize the feature into $H\times W$. For the ray $r_{i}$, its corresponding ray feature can be extracted from $F^{\text{dense}}_{\text{rgb}}$ using ROI alignment [23], with an input size of $8\times 8$ and an output size of $2\times 2$. The ray feature $F_{\text{ray}}$ is obtained by flattening the output of the ROI alignment.
@@
For the voxel feature, early work [13] generates it by sending the position and color of the valid point cloud into a voxel-based PointNet [24, 25]. However, this approach may not fully utilize the semantic information provided by the RGB image. To overcome this limitation, we introduce an early-stage point fusion method. We first encode the geometry embedding of point cloud $P_{\text{xyz}}\in\mathbb{R}^{N\times 16}$ using the relative position of points to their voxel centers. We then project the point cloud with camera intrinsic parameters onto the generated dense color feature map $F^{\text{dense}}_{\text{rgb}}$ and generate point-wise color embedding $P_{\text{rgb}}\in\mathbb{R}^{N\times 16}$ through channel reduction. These two features are concatenated into the fused point feature $P_{\text{fuse}}$. The final voxel feature $F_{\text{voxel}}$ is generated from $P_{\text{fuse}}$ using a voxel feature encoder, which contains a set of voxel-wise maxpooling layers and fully connected layers [26].
@@
Hand Feature Extraction. In addition to the ray and voxel feature, HADR incorporates hand pose as important guidance for depth restoration. One of our insights is that the hand pose can provide both geometric and semantic context, enhancing the accuracy and generalization ability of object reconstruction. As shown in Fig. 4(b), we extract the 21 hand keypoints with a pre-trained hand pose detector. In our experiment, a ResNet-18 [27] is adopted as the backbone to extract the hand feature map from the corrupted depth image, which is then transferred into hand keypoints $X_{\text{hand}}\in\mathbb{R}^{21\times 3}$ through the weighted average regression [28].
@@
Two types of hand features are derived from the detected hand keypoints. We first transfer the keypoints into the hand-wrist coordinate and obtain the absolute hand feature $F_{\text{hand}}^{\text{abs}}$. This feature represents the absolute hand-holding pose, which is invariant to the change of camera position and orientation. The second type of hand feature $F_{\text{hand}}^{\text{rel}}$ is the relative position between each hand keypoint and the center of intersected voxel $v_{j}$. This feature represents the spatial relationship of hand-object interaction. Absolute and relative hand features are flattened and concatenated into the final hand feature $F_{\text{hand}}$.
@@
where $\oplus$ represents the feature-wise concatenation operator, $\gamma(r_{i})$ and $\gamma(d_{ij})$ are the ray direction and voxel position embedding, which has shown to be important for the final accuracy [13]. Two MLPs are used to decode the feature into the predicted terminated position and the existing probability of each ray-voxel pair. We obtain the final predicted depth value by using ray-wise maxpooling [13], as expressed in Equation 1.
@@
III-C H2R Handover Workflow of Transparent Objects
Figure 5: Overview of the proposed handover workflow. The whole handover process is divided into three stages: 1) Wait & Observe, 2) Approach & React, and 3) Grasp & Retrieve.
Based on the proposed hand-aware depth restoration method, we further develop a handover workflow for transparent objects. As shown in Fig. 5, we divide the whole H2R handover process into three stages: 1) Wait & Observe, 2) Approach & React, and 3) Grasp & Retrieve.